OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2m2 + 3 = m double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

\[ 2m^2 -m + 3=0\\ b^2-4 ac = 1 - 24 =-23<0 \] Two imaginary roots

OpenStudy (whpalmer4):

rearrange the equation into the form \[ax^2+bx+c = 0\]the discriminant \((\Delta)\) is then given by \(\Delta = b^2-4ac\) If \(\Delta = 0\) you have a perfect square and a double root if \(\Delta > 0\) you have two real roots if \(\Delta < 0\) you have two imaginary roots If you can't simplify \(\sqrt{b^2-4ac}\) into something without a radical sign, you've got irrational roots.